OpenStudy (anonymous):

Evaluate this exponential expression. (27/8)^4/3

OpenStudy (anonymous):

\[(\frac{3}{2})^{3*\frac{4}{3}} =(\frac{3}{2})^4 = \frac{81}{16}\]

OpenStudy (anonymous):

Can that be simplified at all?

OpenStudy (anonymous):

means u wanna simplify it further its not coz numerator is odd and denominator is even